# Cygnus X-3 as semi-hidden PeVatron

### v1

@@ 5. Discussion
  • Our values for the masses of the binary stars are at the higher end of the range discussed in the literature, and the choice $M_{\rm BH}=5M_{\odot}$ and $M_{\rm WR}=20M_{\odot}$ would be more in line with the analysis of, e.g., Koljonen & Maccarone (2017). This value for $M_{\rm BH}$ would raise the temperature of the accretion disk by only $\sim 30\%$, and we have seen that the disk photons play only a minor role as target. More importantly, the number density of X-ray photons in the state $S_{1}$ would increase by a factor $\simeq 64$, reducing the maximal energy in the $S_{1}$ state by a factor of order ten. Such a reduction of the maximal proton energy would mean that Cygnus X-3 is barely a PeV photon source in the $S_{1}$ state. Alternatively, a two-zone model would have to be considered where the X-ray emission is partly decoupled from the acceleration of hadrons.
  • The accretion rate $\dot{M}$ regulates the total amount of power available, of which the fraction $\eta_{\rm CR}\lambda$ can be channeled into the acceleration of CRs. X-ray observations favour super-Eddington accretion rates and/or a rather massive BH (Veledina et al., 2024). Moreover, radio observations show a very large variability in the radio flux (Gregory et al., 1972; Green et al., 2025). If this variability is connected to variations in $\dot{M}$ or $\eta_{\rm CR}$, then the PeV photon flux should be in an one-zone model correlated with the variation in the radio band. If, on the other hand, a smaller jet-opening angles and larger jet velocities are characteristic for flares (Spencer et al., 2022), then the PeV photon flux could be in the state $S_{2}$ anti-correlated to the radio flares.
  • In the state $S_{2}$, a combined scenario with both DSA at the recolliminaton or termination shock and 2.nd order Fermi throughout the jet or magnetic reconnection close the the BH could be possible. In particular, for a smaller value of $\delta$, the maximal energy achievable in DSA is reduced and this region would not contribute to high-energy photons. However, this region might be still the main contributor to the X-ray emission via leptonic processes. Hence such a case would be a realisation of a two-zone model, where the target density for photo-hadronic interactions is strongly reduced, and rather small BH mases would be unproblematic even for the acceleration via reconnection close to the BH.
@@
  • We have considered a purely hadronic model for the photon emission of Cygnus X-3, without trying to explain self-consistently the X-ray emission by electrons. A combined model where the acceleration of electrons and hadrons is simulated based on same mechanism would help to constrain physical parameters like the magnetic field strength and the size of the acceleration region, but is deferred to a future work.
6. Conclusions

### v2

@@ 5 Discussion
  • Our values for the masses of the binary stars are at the higher end of the range discussed in the literature, and the choice $M_{\rm BH}=5M_{\odot}$ and $M_{\rm WR}=20M_{\odot}$ would be more in line with the analysis of, e.g., Koljonen & Maccarone (2017). This value for $M_{\rm BH}$ would raise the temperature of the accretion disk by only $\sim 30\%$, and we have seen that the disk photons play only a minor role as target. More importantly, the number density of X-ray photons in the state $S_{1}$ would increase by a factor $\simeq 4^{3}=64$, reducing the maximal energy in the $S_{1}$ state by a factor of order ten. Such a reduction of the maximal proton energy would mean that Cygnus X-3 is barely a PeV photon source in the $S_{1}$ state. Alternatively, a two-zone model would have to be considered where the X-ray emission is partly decoupled from the acceleration of hadrons.
  • The accretion rate $\dot{M}$ regulates the total amount of power available, of which the fraction $\eta_{\rm CR}\lambda$ can be channeled into the acceleration of CRs. X-ray observations favour super-Eddington accretion rates and/or a rather massive BH (Veledina et al., 2024). Moreover, radio observations show a very large variability in the radio flux (Gregory et al., 1972; Green et al., 2025). If this variability is connected to variations in $\dot{M}$ or $\eta_{\rm CR}$, then the PeV photon flux should be in an one-zone model correlated with the variation in the radio band. If, on the other hand, a smaller jet-opening angle and larger jet velocities are characteristic for flares (Spencer et al., 2022), then the PeV photon flux could be in the state $S_{2}$ anti-correlated to the radio flares.
  • In the state $S_{2}$, a combined scenario with both DSA at the recolliminaton or termination shock and 2.nd order Fermi throughout the jet or magnetic reconnection close the the BH could be possible. In particular, for a smaller value of $\delta$, the maximal energy achievable in DSA is reduced and this region would not contribute to high-energy photons. However, this region might be still the main contributor to the X-ray emission via leptonic processes. Hence such a case would be a realisation of a two-zone model, where the target density for photo-hadronic interactions is strongly reduced, and rather small BH masses would be unproblematic even for the acceleration via reconnection close to the BH.
@@
  • We have considered a purely hadronic model for the photon emission of Cygnus X-3, without trying to explain self-consistently the X-ray emission by electrons. A combined model where the acceleration of electrons and hadrons is simulated based on the same mechanism would help to constrain physical parameters like the magnetic field strength and the size of the acceleration region, but is deferred to a future work.
6 Conclusions